OpenStudy (anonymous):

half-life of chromium- 51 is 28 days. Starting with an original sample of 12. micrograms, calculate the number of days it will take for the radioactive isotope to drop to .75 micrograms.

OpenStudy (cuanchi):

there is at least two ways to solve the problem one is just trying to figure out if the final amount is a 1/2, 1/4, 1/8, 1/16, 1/32 of the original sample what is going to be 1, 2,3,4,5 half life the other is using the t1/2= 0.69/k equation to calculate the k and then the integrated equation of the first order kinetics to solve the value of t ln[A]=−kt+ln[A]o